### v2

# S4ST: A Strong, Self-transferable, faSt, and Simple Scale Transformation for Data-free Transferable Targeted Attack

Abstract
Transferable Targeted Attacks (TTAs), which aim to deceive black-box models into predicting specific erroneous labels, face significant challenges due to severe overfitting to surrogate models. Although modifying image features to generate robust semantic patterns of the target class is a promising approach, existing methods heavily rely on large-scale additional data and prolonged training, or at least require query access to victim models for method design and tuning. This dependence on prior information undermines the fair evaluation of TTA threats, potentially leading to a false sense of security or unnecessary overreactions. In this paper, we introduce two novel blind estimation measures—surrogate self-alignment and self-transferability—to analyze the effectiveness and correlations of basic transformations, to enhance gradient-based data-free attacks under strict black-box constraints (no access to victim models or their data distributions). Our findings challenge conventional assumptions: (1) Attacking simple scaling transformations uniquely enhances targeted transferability, outperforming other basic transformations and rivaling leading complex methods; (2) Geometric and color transformations exhibit high internal redundancy despite weak inter-category correlations. These insights drive the design and tuning of S4ST (Strong, Self-transferable, faSt, Simple Scale Transformation), which integrates dimensionally consistent scaling, complementary low-redundancy transformations, and block-wise operations. Extensive experiments on the ImageNet-Compatible dataset demonstrate that S4ST achieves a 77.7% average targeted success rate (tSuc) in the 10-Targets (all-source) setting, surpassing existing transformations (+17.2% over H-Aug with only 26% computational time) and state-of-the-art TTA solutions (+6.2% over SASD-WS with 1.2M samples for post-training). Notably, it attains 69.6% and 55.3% average tSuc against three commercial APIs and vision-language models, respectively. This work establishes a new SOTA for TTAs, highlights their potential threats, and calls for a reevaluation of the data dependency in achieving targeted transferability. Codes are available at https://github.com/scenarri/S4ST.
Index Terms:
Adversarial examples, gradient-based attack, transferability, targeted adversarial attack, deep neural networks
1 Introduction
Deep Neural Networks (DNNs) have been shown to be vulnerable to adversarial examples (AEs) [1], which refer to images with small, intentionally crafted perturbations that are usually imperceptible to the human visual system but can fool DNNs. The presence of AEs raises significant concerns when deploying DNNs in security-critical domains, including autonomous driving, face recognition, and medical image analysis. Therefore, finding, understanding, and defending AEs have received significant attention [2, 3, 4, 5]. AEs exhibit an intriguing property known as transferability [6], which means that AEs designed for one DNN can successfully deceive other DNNs with distinct architectures. The transferability of AEs facilitates the implementation of black-box attacks where queries to the victims are forbidden. As a result, significant attention has been paid to enhancing these examples’ transferability in the last decade [7].
While transferable untargeted attacks have been extensively investigated, yielding a substantial body of research [8, 9, 10, 11, 12, 13], Targeted Transferable Attacks (TTAs) are significantly more challenging and consequently less explored. This is primarily due to the exacerbated overfitting of AEs to surrogate models in the context of targeted attacks. More specifically, it is challenging to construct a targeted perturbation that can drive samples to a specific target class within the decision spaces of various unknown models. To achieve effective targeted transferability, directly modifying the original visual features to form robust semantic patterns of the target class has become a mainstream alternative [14, 15], requiring the exploitation of extensive data to model the underlying distribution. To this end, data-reliant TTAs often entail optimizing universal perturbations [16], generators [15, 17, 18], or the surrogate model itself [19, 20, 21]. However, two key drawbacks exist due to their need for tens of thousands of additional data points:
  1. 1.
    The data used to model the distribution is typically the training or testing data of the target model, leading to data leakage and violating the black-box attack setting. This hinders the proper evaluation of threats by the research community.
  2. 2.
    Most of these methods demand substantial time and resources for training. For instance, generation-based methods require training a separate generator for each target class, which limits the further utilization of targeted adversarial examples such as cross-dataset applications and adversarial training.
In this paper, we focus on data-free TTAs, also known as the gradient-based attacks [8, 9], which involve executing gradient descent on any (sample, label) pair, offering both efficiency and flexibility. In the meantime, these are plagued by severe overfitting and poor transferability [22].
The limitations of data-free TTAs have been primarily attributed to the gradient vanishing issue of the cross-entropy loss function in mainstream research, which has spurred the development of a series of alternative loss functions [23, 24, 25, 26]. However, input transformations play a more critical role [10, 27]. These transformations are incorporated into the iterative gradient update process, serving a dual purpose: on one hand, they diversify the image features, and on the other hand, they soften the surrogate’s prediction distribution, thereby mitigating gradient vanishing. This dual mechanism makes the realization of effective data-free TTAs possible. In fact, specialized loss functions lose their efficacy when decoupled from input transformations [24]. Some existing transformations designed for untargeted attacks such as the random scaling & padding transformations in [10] and [28], have demonstrated promising improvements [29, 24]. Nevertheless, there is still a lack of guidelines for designing transformation methods under black-box conditions, and their potential remains largely locked. Therefore, we aim to understand how to design highly effective and efficient transformation methods, to what extent these methods can unlock the potential of data-free TTAs, and whether it is possible to outperform data-reliant methods without access to the target model and its training data.
In this paper, we simplify the above objective to identify a strategy to combine various basic transformations. First, we propose a clever use of the overfitting of targeted AEs. Specifically, the more severe the overfitting, the weaker the impact of AEs on unknown models, indicating a greater misalignment between the surrogate model and unknown models on AEs. Therefore, we directly use the self-alignment of the surrogate model on transformed AEs and clean samples to estimate the alignment towards unknown models on AEs, i.e., the effectiveness of basic transformations. Furthermore, by attacking the surrogate model using various complex transformation methods and computing its self-transferability against basic transformation defenses, we explore the correlations (or redundancies) among basic transformations. Through these two simple yet effective analyses, the following key findings are gained:
  1. 1.
    Simple scaling transformations exhibit outstanding capability in promoting targeted transferability. Compared to other basic transformations, scaling demonstrates a stronger effect on promoting surrogate self-alignment and operates effectively over a wider range of transformation intensities.
  2. 2.
    Geometric and color transformations exhibit high internal similarity but weak inter-category correlation. That is, even if two transformations appear to be less related, there still exists strong redundancy, such as between scaling and rotation transformations. This observation reminds us to cautiously seek complementary enhancements.
Figure 1: Comparison against existing transformation methods at incremental attack iterations. Compared with the previous SOTA transformation, H-Aug [27], S4ST yields an absolute improvement of 14.2% on average tSuc to 83.0%, and a relative reduction of 74.3% on time consumption to 1.45s. Additionally, compared with the previous setting ($T=300$), more attack iterations benefit a lot for potent transformations like our S4ST.
Further, the findings above have enabled the development of a novel and powerful scaling-centered transformation method for data-free TTAs, named Strong, Self-transferable, faSt, and Simple Scale Transformation (S4ST). S4ST amalgamates three efficient strategies to enhance the simple scaling transformation: 1) Modifications that maintain dimensional consistency while injecting greater diversity and reducing computational burden; 2) Integration with complementary transformations that are less correlated with scaling; and 3) Implementation of block-wise scaling. It is worth noting that the proposed S4ST requires no extra data to determine what complementary transformations should be applied and, even for parameter tuning (all done according to the self-transferability analysis in a strict black-box manner)!
In summary, we provide novel insights and methods for designing input transformation methods to enhance TTAs, and we show that under strict black-box conditions—where any access of the black-box model or its data is prohibited—it is possible to complete method design and parameter tuning while achieving SOTA TTA performance. Our work highlights the importance of dedicating resources to the evaluation and defense against targeted AEs in safety-critical applications. Overall, the main contributions of this paper are as follows:
  1. 1.
    We propose two blind estimation measures: surrogate self-alignment and self-transferability against basic transformations. Serving as effective proxies for the black-box transferability of targeted AEs, they enable feasible analyses of the effectiveness and synergistic effect of various basic transformations without accessing any additional data and victim models, significantly diminishing the dependence on empirical or intuitive choices common in the prior art.
  2. 2.
    To the best of our knowledge, we are the pioneers in revealing the unique superior efficacy of the simple scaling transformation in enhancing targeted transferability, and the redundancies within geometric and color transformations. Based on these, we further design S4ST, an advanced scaling-centered transformation that ingeniously integrates modified scaling with complementary transformations and leverages the benefits of block-wise operations, adhering to a strict black-box manner.
  3. 3.
    Extensive and comprehensive experiments on the ImageNet-Compatible dataset substantiate the effectiveness and efficiency of S4ST. It not only outperforms existing transformation techniques in terms of both effectiveness and efficiency (see Fig. 1) but also surpasses SOTA TTA solutions that depend on 50k to 1.2 million training samples (see Table VII), all without requiring additional data. It also exhibits exceptional transferability to commercial APIs and vision-language models. It is found to significantly improve the exploitation on existing visual elements to produce highly transferable target semantics, circumventing the necessity to learn this capability from extensive multi-class data.
The rest of this paper is organized as follows. Section 2 introduces the problem definition, related works, and experimental validations of existing consensuses in designing transformation methods. Section 3 elaborates on the motivation, analysis, findings, and details of our method. Section 4 offers the experimental settings, results, and analysis. Finally, Section 5 concludes this work.
2 Background and Preliminary
2.1 Problem Setting and Baseline Approach
Consider an image classifier $F$ that predicts a label among $K$ categories, $\bm{y}=\{y_{i}|y_{i}\in\mathbb{Z},0\leq i<K\}$, for an input with RGB channels, $\bm{x}\in\mathbb{R}^{3\times H\times W}$, where $H$ and $W$ denote the image height and width, respectively. Without loss of generality, we denote $F=C\circ f$, where $f(\bm{x})=\bm{p}\in[0,1]^{K}$ represents the prediction distribution and $C(\bm{p})=y_{\operatorname{argmax}(\bm{p})}$ indexes the label with the highest probability. TTA aims to manipulate black-box classifiers, $G=C\circ g$, to produce an arbitrary targeted prediction, $y_{t}$, for any input:
$\centering G(\bm{x}^{adv})=y_{t},\quad\text{s.t.}\quad\|\bm{x}^{adv}-\bm{x}\|_{p}\leq\epsilon,\@add@centering$ (1)
where $\bm{x}^{adv}$ is the AE crafted to be within a $p$-norm ball centered at $\bm{x}$ with radius $\epsilon$ (the perturbation budget). In the context of transfer attacks, $\bm{x}^{adv}$ is generated using the surrogate model $f$:
$\centering\bm{x}^{adv}=\underset{\bm{x}^{\prime}:\left\|\bm{x}^{\prime}-\bm{x}\right\|_{p}\leq\epsilon}{\arg\min}\mathcal{L}\left(f(\bm{x}^{\prime}),y_{t}\right),\@add@centering$ (2)
where $\mathcal{L}$ is the loss function designed to induce the targeted misclassification. In line with the established protocol [24], we employ the gradient-based method, TMI [9, 8], as the baseline for date-free attacks:
$\bm{x}^{adv}_{0}=\bm{x},\quad\bm{g}_{0}=\bm{0},$ (3)
$\bm{g}_{i+1}=\mu\cdot\bm{g}_{i}$ $+\frac{\bm{W}*\nabla_{\bm{x}}\mathcal{L}(f(\mathcal{T}(\bm{x}^{adv}_{i}),y_{t}))}{\left\|\bm{W}*\nabla_{\bm{x}}\mathcal{L}(f(\mathcal{T}(\bm{x}^{adv}_{i}),y_{t}))\right\|_{1}},$
$\bm{x}^{adv}_{i+1}$ $=\Pi_{\epsilon}(\bm{x}^{adv}_{i}-\alpha\cdot\operatorname{sign}(\bm{g}_{i+1})).$
Here, The step size $\alpha$ and perturbation budget $\epsilon$ are configured to 2/255 and 16/255, respectively. The momentum accumulation decay factor $\mu$ is set to 1 [8], and a 5$\times$5 Gaussian kernel [9] $\bm{W}$ is used for TI. We adopt the margin-based cross-entropy loss [25] and the attack iteration is extended to $T=900$ (compared to 300 in previous work) in keeping with the increased diversity afforded by the most advanced transformation methods $\mathcal{T}$.
2.2 Transferable Targeted Attacks
In this paper, current solutions for TTA are categorized into data-free and -reliant approaches as elaborated upon in the following.
2.2.1 Data-free TTAs
Data-free TTAs employ gradient-based iterative optimization on a target (sample, label) pair using a surrogate model [30, 8, 9, 2]. While effective in untargeted attacks, the cross-entropy loss faces challenges in TTAs due to gradient vanishing issues [23, 24]. Let $\bm{\phi}$ denote the representation produced by $f$, such as features preceding the final linear layer. The logit values are computed as $l_{i}=w_{i}^{\top}\bm{\phi}+b_{i}$, with $w_{i}$ and $b_{i}$ being the class-specific weight and bias, respectively. The normalized probability with respect to each category is given by the SoftMax function: $p_{i}=\frac{e^{l_{i}}}{\sum e^{l_{j}}}$. The gradient
$\centering\frac{\partial\mathcal{L}_{\text{CE}}}{\partial\bm{x}}=\frac{\partial\mathcal{L}_{\text{CE}}}{\partial\bm{\phi}}\frac{\partial\bm{\phi}}{\partial\bm{x}}=[\left(p_{t}-1\right)w_{t}^{\top}+\sum_{i\neq t}p_{i}w_{i}^{\top}]\frac{\partial\bm{\phi}}{\partial\bm{x}}\@add@centering$ (4)
tends to vanish as the model becomes confident ($p_{i\neq t}\rightarrow 0$). To address this, Li et al. introduced a Poincaré distance-based triplet [23] loss to facilitate effective optimization. The Logit loss [24] proposed by Zhao et al. directly maximizes the target logit and provides a direct solution to the gradient vanishing dilemma. Subsequent studies have explored ways to soften $\mathcal{L}_{\text{CE}}$ to stabilize the optimization process, such as through margin-based logit calibration [25] or an adversarial optimization scheme [26].
Beyond the design of loss functions, it is the image transformations, as elaborated upon below and in Sec. 2.3, that have proven crucial for achieving effective data-free TTAs and have so far been underexplored. Zhao et al. [24] demonstrated that the Diverse Inputs (DI) transformation [10] significantly enhances the transferability of targeted attacks, demonstrating that AEs rarely transfer successfully without it. An exception can be found in [29] where targeted transferability is boosted via mixing the clean and adversarial features at intermediate DNN layers. However, it also largely depends on transformation techniques. Although we categorize the aforementioned methods as data-free, it is important to emphasize that they still require leveraging feedback from black-box models to tune their parameters. The optimization process was observed to require substantially more iterations (approximately 300) to converge. Since then, the DI method has become a standard benchmark in subsequent research. However, efforts to understand which transformations are essential for targeted transferability remain scarce. Wei et al. [27] provided empirical evidence suggesting that DI-transformed images diversify the discrimination regions of the surrogate model across iterations, thereby expanding the final discrimination region with respect to the targeted label in black-box models. They also introduced the T/H-Aug method, which applies random combinations [31] of attention-deviation transformations. Byun et al. proposed the Object-based DI (ODI) [32] transformation, inspired by the intuition that AEs transferable across various 3D shapes and viewpoints could more effectively compromise black-box models. In [33], the authors add an extra inference branch to pursue a consistent adversarial effect between random crops and the whole image.
In this work, we explore in greater depth the role of image transformations within data-free TTAs. We also challenge the limitations of existing empirical insights and intuitive approaches, proposing new methods to analyze various basic transformations. Building on these, we have developed the S4ST transformation, which has led to notable performance improvements in a strict black-box setting.
2.2.2 Data-reliant TTAs
Data-reliant TTAs seek to improve targeted transferability by leveraging the data distribution through a substantial volume of additional data. Targeted universal attacks [14, 16] develop a single perturbation across a large dataset, enabling the universal attack of multiple unseen samples and demonstrating cross-model transferability. Inkawhich et al. chose to train auxiliary models at intermediate layers to capture and attack the feature distribution relative to the target class [34]. Springer et al. reveal that adversarially training surrogate model against small-magnitude AEs yields substantial improvement in targeted transferability [35]. The score-matching techniques are borrowed by Intrinsic Adversarial Attack (IAA) [19] and Distribution-Relevant Attack (DRA) [20] to respectively tune the surrogate models’ architectural parameters and weights for modeling the data distribution. Wu et al. propose the Sharpness-Aware Self-Distillation and Weight Scaling (SASD-WS) method, which combines the sharpness-aware optimization [36] and model distillation for model self-enhancement and results in better transferability.
The pioneering generative framework for TTA, known as Transferable Targeted Perturbation (TTP) [15], involves training a generator on extensive data to align with a specified target class in terms of both global distribution and local neighborhood structure. Zhao et al. introduced the Minimizing Maximum Model Discrepancy (M3D) attack [18], which discriminates and targets two surrogate models (the original and its duplicate). Conditional Generative Semantic Pattern (C-GSP) [37] enables perturbation generators to attack multiple labels simultaneously by incorporating conditional learning.
Although these methods are effective and efficient in inference, their practicality is constrained by the requirement for extensive training data from the same domain or even the same dataset as the black-box models and the laborious retraining process for large amounts of target labels. By building on our proposed S4ST transformation, data-free TTAs have surprisingly surpassed SOTA data-reliant techniques, all while preserving remarkable flexibility and effectiveness.
2.3 Input Transformations for Transferability
Image transformation is an effective and necessary strategy in training deep vision models, effectively mitigating the need for extremely large quantities of data and annotations [38, 39, 40]. For example, horizontal flips and random resized cropping are commonly used in natural image understanding tasks, significantly boosting generalization performance. Furthermore, transformation strategies are typically data-dependent and can be conveniently implemented in a plug-and-play manner across different model architectures. The same benefits extend to the domain of adversarial attacks. There are two primary streams of transformation strategies for enhancing adversarial transferability: 1) Dynamically searching for optimal combinations from a vast pool of transformations during the attack iteration [41, 42, 43]; and 2) Consistently applying a predefined transformation in each iteration. In this paper, we focus exclusively on the latter to avoid the need to access black-box models for optimization. In addition to TTA-oriented transformation methods [27, 32], we also examine several prominent works in the field of untargeted attacks. This is to validate whether the prevalent consensuses in transformation design are applicable in the context of TTA. An overview of the methods involved is presented in Fig. 2, introduced below.
Xie et al. first introduced the DI method [10], which involves random resizing and padding, and has shown a significant boost in transferability. The Resized DI (RDI) approach [28] rescales the DI-transformed image back to the original size, creating a reduced-scale version. The Scale-Invariant (SI) technique [44] computes the average gradient with respect to images scaled linearly by powers of two, repeated five times. Admix [45] blends the original image with those from other classes. The Spectrum Simulation Attack (SSA) [46] introduces noise within the frequency domain. The Structure Invariant Attack (SIA) [11] and Block Shuffle & Rotation (BSR) [12] employ random transformations and shuffling & rotation operations on image blocks, respectively. The Deformation-Constrained Warping (DeCoWa) method [13] applies elastic deformation coupled with an internal min-max optimization step.
Figure 2: Illustration of the original image and its transformed versions.
Figure 3: Scatter diagrams depicting relationships between black-box transferability and (a) diversity, (b) attention deviation, and (c) gradient magnitude. As indicated by the Pearson Correlation Coefficients (PCCs) in the top left corner of each figure, all three metrics exhibit weak correlation to transferability in the cross-method case, impeding their application in identifying what transformations are effective and designing a complex transformation from scratch.
2.4 Failures of Existing Consensuses
We scrutinize the prevailing consensuses commonly referenced in the development of transformation methods for transfer attacks. Our objective is to ascertain their robustness as metrics reflecting targeted transferability and to determine their utility in guiding the design of transformations for TTAs. To achieve this, we assess the black-box transferability of various transformation techniques under both default and expanded hyperparameter configurations. Here, the black-box transferability is calculated by the variance in predicted probabilities of black-box models with respect to the target class $y_{t}$ between the adversarial (generated by $\mathcal{T}$-TMI) and clean samples:
$\centering\textit{T}_{\mathcal{T}}=\mathbb{E}_{\bm{x}\in\mathcal{X},g\in\mathcal{G}}\left(g(\bm{x}_{\mathcal{T}}^{adv},t)-g(\bm{x},t)\right).\@add@centering$ (5)
We then examine the presence of a positive (or linear) correlation between their transferability and the quantitative measures derived from the established consensus. The consensuses under investigation and their corresponding measures are as follows:
  • Diversity. Initially, researchers emphasized the need for transformations that preserve loss [45, 10, 44, 9], ensuring that the surrogate model’s output remains relatively constant post-transformation. The intensity of transformations was thus limited to a narrow scope to maintain this characteristic. However, more recent advancements have shifted towards utilizing gradient averaging over multiple transformed samples (20$\times$) to achieve stable gradient updates [12, 11, 13, 46]. This shift implies that increased diversity might lead to improved transferability. We quantify diversity by calculating the average cross-entropy loss values of the surrogate model across 50,000 images from the ImageNet-1K validation set.
  • Attention Deviation. Dong et al. first pointed out that discrepancies in discriminative regions between models and defensive strategies can hinder adversarial transferability [9]. They advocated for the use of translation transformations to extend the impact of adversarial perturbations to wider attention regions. Following this, other studies have also utilized attention as a universal objective for crafting transferable attacks [47, 48]. More recently, BSR [12] and H-Aug [27] have set the new SOTAs for untargeted and targeted attacks by disrupting attention through image block shuffling and random attention-deviation transformations, respectively. We adopt the attention deviation metric proposed by Wei et al. [27], which computes the Intersection over Union (IoU) of Grad-CAM [49] activations concerning the target label between original and transformed images. The results are averaged over 1,000 images with distinct target labels.
  • Gradient Magnitude. A significant portion of research on data-free TTAs has been devoted to addressing the issue of gradient vanishing, particularly in studies focusing on loss functions [24, 33, 26, 25]. The magnitude of gradients has often been used as an indicator to gauge the efficacy of different approaches. It is hypothesized that larger gradient magnitudes could lead to more efficient updates of perturbations. To test this hypothesis, we examine the gradient magnitudes produced by various transformations under an identical loss function. Specifically, the gradient magnitude is measured by the mean $\ell_{2}$-norm of gradients over 900 $\mathcal{T}$-TMI iterations and is averaged across 1,000 images with distinct target labels.
The results are presented in Fig. 3. Within the intra-method analyses—where the relationship between different hyperparameter configurations of individual methods is evaluated—a positive or linear relationship may be discernible. Notably, the expanded parameters for the DI method expose the constraints of experimental setups in previous research. Nevertheless, this only suggests that these metrics can provide empirical guidance for parameter scaling within specific transformations. Importantly, there is a significant confusion of these metrics across different methods, which complicates the task of developing new transformation methods from the ground up.
3 The Proposed Method
This section systematically delineates the proposed method. It integrates the motivation derived from previous analyses, delineates two plausible strategies for gauging the impact of transformation methods on targeted transferability, and culminates in the introduction of a powerful yet efficient transformation technique to augment data-free TTAs.
3.1 Motivation
In our research, we fervently promote the investigation into robust blind estimators for transformation-induced targeted transferability. To this end, we aspire to establish design principles for formulating transformation techniques. Our inquiry specifically targets questions such as which transformations prove most effective and how to determine the synergistic effects of combining various transformations. We anticipate that our findings will identify transformations that facilitate highly effective data-free TTAs. Put succinctly, our goal is to surpass the efficacy of data-reliant methods while preserving efficiency and the ability to handle arbitrary (sample, label) pairs. This is accomplished without resorting to additional data or retraining models, thereby striving for an optimal balance between performance and practicality. Moreover, the estimators in question should conform to the prerequisites of black-box transfer attacks, thereby ensuring their applicability to a broader spectrum of data beyond natural imagery. To begin, we focus on basic transformations and subsequently investigate their potential synergistic enhancements. The specifics of the basic transformations evaluated and their respective intensities $\mathcal{S}$ are meticulously defined and expounded upon in Appendix A.
3.2 Assessing Effectiveness via Surrogate Self-Alignment
First, we manage to assess the effectiveness of attacking basic transformations to enhance the targeted transferability. This investigation is challenging due to the prohibition of direct evaluations on black-box models. To circumvent this, we suggest an approach rooted in the concept of model alignment. The underlying rationale is straightforward: models trained on similar distributions exhibit a degree of inherent alignment. AEs that capitalize on this alignment between the surrogate and black-box models are more likely to transfer. However, the persisting discrepancies [50] between models, along with the overfitting of gradient-based optimization, sharply misalign the surrogate model and black-box models on AEs. In Fig. 4, we illustrate the average alignment metric between the surrogate and black-box models (defined subsequently in Eq. (6)) across the attack iterations. It becomes apparent that direct utilization of the surrogate model’s gradients can lead to substantial misalignment with black-box models, culminating in poor transferability. Although transformations can mitigate this misalignment, it prompts the question: how can one blindly assess the efficacy of these basic transformations in facilitating this mitigation?
Figure 4: Alignment between the surrogate model and black-box models on targeted AEs, which reflects the black-box transferability resulted by different transformations (see Table IV). The vanilla gradient update, i.e., TMI, quickly misaligns the surrogate from black-box models and leads to poor transferability.
Our solution leverages what we term surrogate self-alignment induced by basic transformations. To simplify our analysis, we first examine the similarity between representations of individual samples. An illustration is depicted in Fig. 5. Given the inherent alignment between models $f$ and $g$, we can reasonably postulate that the cosine similarity of their representations, $\text{CosSim}(\bm{\phi}(\bm{x}),\bm{\psi}(\bm{x}))$, is significantly greater than zero. More informally, we may suggest that $\bm{\phi}(\bm{x})\approx\bm{\psi}(\bm{x})$. For an AE $\bm{x}^{adv}$, crafted using a weakly-transferable algorithm like TMI that has a limited impact on the black-box model $g$, we hypothesize that $\bm{\psi}(\bm{x}^{adv})\approx\bm{\psi}(\bm{x})\approx\bm{\phi}(\bm{x})$, with all three representations being markedly different from $\bm{\phi}(\bm{x}^{adv})$. This difference becomes more pronounced with each iteration of the attack. Image transformations can enhance attack effectiveness by harnessing the mutual alignment between $f$ and $g$, with the aim of making $\bm{\psi}(\bm{x}^{adv})$ closer to $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))$. In line with the principles of black-box attacks, we suggest an extension of this strategy such that $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))\approx\bm{\phi}(\bm{x})$ (self-alignment).
Figure 5: The colored numbers indicate the cosine similarity between representations generated by the surrogate model (ResNet-50) and the black-box model (ConvNeXt), where the AE $\bm{x}^{adv}$ is created based on the TMI algorithm. On the clean sample, these two models exhibit a high degree of alignment. However, AEs with weak transferability have limited impact on the black-box model, which suggests that we can introduce an image transformation $\mathcal{T}$ to make $\bm{\phi}(\mathcal{T}(\bm{x}^{adv}))$ more closely approximate $\bm{\phi}(\bm{x})$ (self-alignment), thereby enhancing the approximation to $\bm{\psi}(\bm{x}^{adv})$ (alignment) and improving transferability.
We then expand the concept of similarity between individual representations to encompass the mutual alignment across the entire dataset and multiple black-box models. Let $\Phi=\{\bm{\phi}(\bm{x}_{i})\}_{i=1}^{b}$, $\Phi_{\widehat{\mathcal{T}}_{s}}=\{\bm{\phi}(\widehat{\mathcal{T}}_{s}(\bm{x}_{i}))\}_{i=1}^{b}$, and $\Psi=\{\bm{\psi}(\bm{x}_{i})\}_{i=1}^{b}$ denote the sets of representations produced by models $f$, $f\circ\widehat{\mathcal{T}}_{s}$, and $g$ for a set of samples $\mathcal{X}=\{\bm{x}_{i}\}_{i=1}^{b}$, respectively. And $s$ represents the intensity level of a basic transformation $\widehat{\mathcal{T}}$. The superscript $(\cdot)^{adv}$ indicates the corresponding representations for the adversarial counterparts of $\mathcal{X}$, which are crafted using the TMI algorithm and model $f$. To quantify the mutual alignment between two sets of representations, such as $\Phi$ and $\Psi$, we measure the overlap of their respective $k$-nearest neighbors [51] throughout the dataset. This is formalized as:
$\operatorname{Alignment}\left(\Phi,\Psi\right)=$ (6)
$\mathbb{E}_{g\in\mathcal{G},(\bm{\phi},\bm{\psi})\in(\Phi,\Psi)}$ $\frac{1}{k}\left|d_{\text{knn}}(\bm{\phi},\Phi\setminus\bm{\phi})\cap d_{\text{knn}}(\bm{\psi},\Psi\setminus\bm{\psi})\right|,$
where $d_{\text{knn}}$ returns the set of indices corresponding to its $k$-nearest neighbors, with $k$ set to 100 for a dataset comprising $b=1,000$ images.
TABLE I: Mutual alignment between surrogate model and multiple black-box models.
Case Alignment
$\operatorname{Alignment}(\Phi,\Psi)$ 0.42
$\operatorname{Alignment}(\Psi,\Psi^{adv})$ 0.32
$\operatorname{Alignment}(\Psi^{adv},\Phi^{adv})$ 0.12
$\operatorname{Alignment}(\Phi^{adv},\Phi)$ 0.10
Figure 6: The metrics of self-alignment and alignment exhibit a high degree of similarity. Notably, the scaling transformation demonstrates distinctive effectiveness, simultaneously inducing high alignment strength and a large effective range of transformation intensities.
Table I presents the alignment metrics between models for various combinations. The results are consistent with those illustrated in Fig. 5, demonstrating that the alignment scores for $\operatorname{Alignment}(\Phi,\Psi)$ and $\operatorname{Alignment}(\Psi,\Psi^{adv})$ are not only comparable but also significantly exceed those of the other two scenarios. This substantiates the efficacy of our motivation, leveraging the self-alignment, $\operatorname{Alignment}(\Phi^{adv}_{\widehat{\mathcal{T}}_{s}},\Phi)$, as a proxy for desired alignment, $\operatorname{Alignment}(\Phi^{adv}_{\widehat{\mathcal{T}}_{s}},\Psi^{adv})$, to blindly gauge the impact of different transformations on the transferability of attacks.
Fig. 6(a) delineates the self-alignment metric for the basic transformations under study at varying intensities. Notably, scaling not only yields a significantly enhanced surrogate self-alignment compared to other geometric transformations but also demonstrates a wider effective intensity range, highlighting its potential as a key transformation for our strategy. To corroborate these preliminary assessments and findings, we conduct comparative experiments among these basic transformations in Section 4.2.2.
Effectiveness of the surrogate self-alignment metric. We further validate the effectiveness of surrogate self-alignment in blindly estimating the real alignment towards unknown black-box alignment. Please see Figs. Fig. 6(a) and Fig. 6(b), these two metrics closely mirror each other, reflecting both the individual transformation trends at distinct intensities and the comparative magnitudes across transformations, thereby corroborating our analysis.
3.3 Revealing Correlation via Self-Transferability
We further investigate the correlations among different basic transformations, which are also achieved without utilizing the target model or its training data. Specifically, we first generate targeted AEs using diverse complex transformation methods (as described in Sec. 2.4). Then, we perform PCC tests on their self-transferability when attacking the surrogate model combined with various basic transformations. This allows us to obtain the correlation relationships between the feature diversity induced by different basic transformations. The underlying rationale is to treat basic transformations as a defensive layer [52] of the surrogate model and use their empirical expectations under different intensities as an estimation of black-box models. If an AE exhibits strong transferability against black-box models, it should also exhibit strong self-transferability against the surrogate model under basic transformation defenses. Furthermore, if the self-transferability against two basic transformations shows a strong correlation, it indicates redundancy between them. Through this analysis, we aim to provide a foundation for the combination and redundancy reduction of our method design. The self-transferability is defined as the mean efficacy of targeted AEs across varying transformation intensities:
$\textit{ST}_{\widehat{\mathcal{T}}}=\mathbb{E}_{\bm{x}\in\mathcal{X},s\in\mathcal{S}}\left(f\circ\widehat{\mathcal{T}}_{s}(\bm{x}_{\mathcal{T}}^{adv},t)-f\circ\widehat{\mathcal{T}}_{s}(\bm{x},t)\right).$ (7)
Figure 7: PCCs between the self-transferability against 12 basic transformations, based on AEs obtained from 12 complex transformations under a total of 40 hyperparameter configurations.
The results of PCC tests for all studied basic transformations are detailed in Fig. 7. It reveals that there are strong internal correlations within geometric transformations and within color transformations, but the correlation between these two categories is relatively weak. Given this, it is more strategic to target a few highly effective basic transformations to optimize overall performance, rather than indiscriminately combining overlapping operations. For example, considering the significant utility of the scaling transformation we previously reported, it may be more advantageous to combine it with color transformations rather than additional geometric transformations to achieve a synergistic enhancement. And the proposed self-transferability measure can assist us in conducting clear and rigorous analyses under strict black-box conditions, see Sec. 3.4.3 for details.
Figure 8: Scatter diagrams depicting relationships between black-box transferability and self-transferability to (a) rotation and (b) scaling transformations.
Effectiveness of the self-transferability measure. Here we validate whether the proposed self-transferability measure, i.e., Eq. (7), serves as a high-quality estimation of the black-box transferability, i.e., Eq. (5). In Fig. 8, we present scatter plots for rotation and scaling transformations, revealing notably tighter correlations in the cross-method scenario compared to previous metrics (as in Fig. 3). The self-transferability of AEs against these basic transformations is observed as an effective reflection of the trends observed in their black-box transferability, applicable to both cross-method and intra-method contexts. The PCCs for all studied basic transformations are detailed in Fig. 9, from which we can discern the following:
  1. 1.
    The AEs’ self-transferability against most geometric transformations demonstrates a consistently high-quality linear correlation with their black-box transferability. For example, attacking features at varying frequency-domain distortions (which often induce color shifts) [46], elastic deformations [13], and random block-wise transformations [11]—despite their seemingly disparate natures—uniformly impact the exploitation of features across scales. This correlation persists even with more complex transformations and a range of geometric ones, such as rotation, shear, and translation.
  2. 2.
    The self-transferability to color transformations shows a weak correlation with black-box transferability in cross-method comparisons, but a robust correlation in the intra-method cases of SSA and SIA. This can be ascribed to the fact that both transformations pertain to the manipulation of image color.
Figure 9: PCCs calculated between the black-box transferability against 14 black-box models and the self-transferability against various basic transformations, based on AEs obtained from 12 existing attacks.
The above results demonstrate that the self-transferability against basic geometric transformations, such as rotation and scaling, can serve as an excellent estimator for black-box transferability. This makes it possible to design transformation methods and perform parameter tuning under strict black-box conditions. We leave that for future study.
3.${}^{\textbf{4}}$
Drawing on the insights and findings from our research, we introduce a novel transformation methodology focused on scaling, named Strong, Self-transferable, Fast, and Simple Scale Transformation (S4ST), designed to enhance data-free TTAs. Fig. 10 illustrates the schematic overview of S4ST, with detailed explanations to follow.
Figure 10: The proposed S4ST encompasses the Base operation, which is based on scaling, complemented by two orthogonal enhancements—Aug and Block.
3.4.1 Overview
The essence of the S4ST framework is founded on the basic scaling transformation, informed by our preceding analysis, and seeks synergistic benefits with complementary basic operations. In alignment with recent advancements, our framework also integrates strategies for image block manipulation. To accommodate block-wise scaling and alleviate the computational demands associated with upscaled inputs, we have refined the scaling transformation to preserve dimensional consistency. As a result, the S4ST framework is segmented into three distinct components: S4ST-Base, S4ST-Aug, and S4ST-Block, each reflecting one of the three primary motivations identified.
3.4.2 S4ST-Base
We define the intensity of the S4ST-Base with the variable $r$, which is distinct from the definition of $s$ in Appendix A. Consider a sample $\bm{x}$ with dimensions $H\times W$ and an intensity $r>1$. With a probability $p_{r}$, S4ST-Base selects a relative scale factor $r^{\prime}$ from a uniform distribution $\mathcal{U}(\frac{1}{r},r)$. If $r^{\prime}<1$, the sample $\bm{x}$ is downscaled to dimensions $r^{\prime}H\times r^{\prime}W$, followed by random zero-padding to restore its original size. In contrast, if $r^{\prime}>1$, a square patch of dimensions $\frac{1}{r^{\prime}}H\times\frac{1}{r^{\prime}}W$ is randomly cropped from the image and then upscaled to the initial dimensions.
Scaling transformations are also utilized in existing approaches, such as DI [10] and RDI [28]. The innovations of our work compared to these methods are threefold: 1) Our black-box condition analysis singles out scaling as the foremost effective basic transformation other than empirical choice; 2) We ascertain that an extended range of scale transformations can lead to further enhancements; 3) We introduce a versatile scaling transformation that includes both enlargement and reduction, unlike previous methods that typically focus on one direction. These advancements enable our S4ST-Base to stand out for its combined effectiveness and efficiency. A comprehensive comparison among variations of the scaling transformation is presented in Section 4.2.3.
3.${}^{\textit{4}}$-Aug
Figure 11: PCCs between the self-transferability to scaling and other basic transformations, derived from 11 intensity levels using S4ST-Base. This evaluation did not involve any black-box models.
Guided by the insights gained from Fig. 9, we aimed to integrate a suite of complementary basic transformations into the S4ST-Base to further boost its efficacy. With a probability $p_{\textit{aug}}$, S4ST-Aug initially applies a randomly chosen transformation from an equitable selection pool of basic transformations to the entire input image. To compile this pool, we execute S4ST-Base across a spectrum of intensities $r\in[1.02,1.04,1.06,1.08,1.1,1.2,1.4,1.6,1.8,2.0,2.2]$. We then evaluate the linear correlation of self-transferability for this array of AEs over a range of basic transformations, with the results displayed in Fig. 11. We select five basic transformations that demonstrate the lowest PCCs with $\textit{ST}_{\text{Scaling}}$. Additionally, we omit the translate and crop transformations due to their significant overlap with S4ST-Base. Following the parameterization in torchvision [^1] https://github.com/pytorch/vision/tree/main/torchvision/transforms, the transformation pool for S4ST-Aug includes the following basic transformations with their empirically determined parameters:
  • Flip randomly flips the image vertically or horizontally.
  • Brightness/Contrast/Saturation randomly alters the brightness/contrast/saturation factor from 1 to $s^{\prime}\sim\mathcal{U}(0,2)$.
  • Hue randomly adjusts the hue factor from 0 to $s^{\prime}\sim\mathcal{U}(-0.5,0.5)$.
3.4.${}^{\textit{4}}$-Block
Recent research has underscored the value of block-wise transformations in enhancing adversarial transferability. With S4ST-Block, we advance this concept by randomly dividing the image into $m$ blocks and applying S4ST-Base to each block, where $m=m_{h}\times m_{w}$ indicates partitioning the image into $m_{h}$ and $m_{w}$ segments along the coordinate axes (not specifying horizontal or vertical orientation in order).
We acknowledge that the design of the block-wise transformation is heavily inspired by prior works such as SIA [11] and BSR [12]. It is important to reiterate that our design is rooted in the analysis conducted within this study, as opposed to being purely based on intuition or empirical choice, in accordance with the shortcomings of existing paradigms discussed in Section 2.4. This analytical approach yields notable differences between our proposed method and pre-existing ones: 1) Our method solely concentrates on block-wise scaling, deduced from our self-alignment estimation concerning the efficacy of various basic transformations, unlike SIA’s arbitrary transformations for each image block or BSR’s shuffling and rotation operations; 2) The selection of complementary transformations in our approach is guided by correlation analyses of self-transferability under black-box scenarios (as depicted in Fig. 11), rather than the intuitive selections employed by SIA and BSR. Comparisons against these methods are provided in Section 4.3.
3.4.5 Black-box Parameter Selection Strategy
Unlike most existing methods that rely on accessing black-box models for hyperparameters tuning, the parameter configuration of the proposed S4ST method also adheres to the strict black-box setting. First, our parameter optimization objective is to maximize the average self-transferability across the 12 studied basic transformations at different intensities, rather than black-box transferability involving the victim models. Second, we employ Bayesian optimization to simultaneously search for the four parameters of S4ST: $[p_{\textit{r}},r,p_{\textit{aug}},m]$, avoiding suboptimal results from optimizing individual components separately. The optimization is performed on a randomly selected subset of the ImageNet-Compatible dataset (50 images), consisting of 10 initial random explorations and a total of 100 trials. The final resulting setting is $p_{\textit{r}}=0.9,r=1.7,p_{\textit{aug}}=1.0,m=6$.
4 Experiments
In this section, we detail the experimental setup, and scrutinize the analysis and the ablation of components for our proposed method, showcasing its novel aspects in contrast to existing techniques. This is followed by extensive experiments designed to compare our method with the existing transformation techniques and SOTA approaches for TTAs.
4.1 Experimental Settings
Dataset. Our experiments are conducted on the ImageNet-Compatible dataset [^2] https://github.com/cleverhans-lab/cleverhans/tree/master/cleverhans_v3.1.0/examples/nips17_adversarial_competition/dataset, a benchmark commonly utilized for evaluating TTA performance [24, 33, 32]. Released for the NIPS 2017 adversarial competition, it encompasses 1,000 images with dimensions of 299$\times$299, each paired with a targeted label. For our evaluated models, we resize the images to 224$\times$224.
Models. In line with prior studies, we predominantly generate AEs using ResNet-50 [53] and transfer them to a variety of dissimilar victim models. These include CNN architectures—{MobileNet-v2 (MNv2) [54], EfficientNet-b0 (EN) [55], ConvNeXt-small (CNX) [56], Inception-v3 (INv3) [57], Inception-v4 (INv4) [58], Inception ResNet-v2 (IRv2), and Xception (Xcep) [59]} and vision transformers—{ViT [60], SwinT [61], MaxViT [62], Twins [63], PiT [64], TNT [65], and DeiT [66]}. We source their pre-trained weights from the torchvision [^3] https://github.com/pytorch/vision/tree/master/torchvision/models and timm [67] libraries. Supplemental black-box transfer experiments are performed on defense models and real-world APIs, with further details provided in subsequent sections.
Competitors. As for existing transformation techniques, we compare our S4ST with DI [10], RDI [28], SI [44], Admix [45], SSA [46], ODI [32], SIA [11], BSR [12], T-Aug [27], and H-Aug [27]. We also compare S4ST-TMI with SOTA solutions including SASD-WS-DI-TMI [21], DRA-DI-TMI [20], IAA [19], Clean Feature Mixup (CFM-RDI-TMI) [29], Self-Universality (SU-DI-TMI) [33], Little Robust Surrogate (LRS-DI-TMI) [35], Dominant Feature Attack (DFA) [14], TTP [15], M3D [18], and C-GSP [37]. We adhere to the implementations detailed in their respective publications and the TransferAttack [^4] https://github.com/Trustworthy-AI-Group/TransferAttack code base.
Metrics. We gauge the efficacy of different methods by their average tSuc, i.e., the proportion of successful transfers across the dataset from the surrogate to the victim models. Additionally, we evaluate the time required by different transformations to craft AEs. All experiments are conducted on an NVIDIA GeForce RTX 4090 GPU using the PyTorch [68] framework. For consistency, the batch size is set to 10 across all experiments, ensuring that no method surpasses the memory threshold, thereby maintaining fairness in the comparison of time consumption metrics.
[…]
4.2.1 Component Ablation
We first detail the ablation outcomes for the components of S4ST with previously reported configuration, as illustrated in Tables II and III. Remarkably, S4ST-Aug and S4ST-Block exhibit satisfying synergy and collectively contribute to a notable increase in performance, achieving an 83.0% average tSuc. Additionally, the pool for S4ST-Aug, which consists of five transformations, demonstrates a degree of complementarity and robustness, with the exclusion of any single transformation having a minimal impact on the overall performance.
TABLE II: Results by different combinations of the S4ST components.
S4ST-Base S4ST-Aug S4ST-Block $[p_{\textit{r}},r,p_{\textit{aug}},m]$ tSuc Time
[…]
TABLE III: Results by excluding complementary augmentations from S4ST.
S4ST -Flip -Contrast -Brightness -Saturation -Hue
83.0 80.8 82.6 82.7 82.7 82.6
4.2.2 Scaling versus Other Basic Transformations
Here we validate the selection rationale for the scaling transformation, evaluating the judiciousness of transformation choices through the lens of the proposed self-alignment evaluation. To this end, we generate adversarial samples employing each basic transformation under consideration. With a probability of 0.9, given a transformation strength $s$, each transformation is executed with an intensity $s^{\prime}\in\mathcal{U}(0,s)$, selecting a random variant for the surrogate model to process and compute gradients. The outcomes illustrated in Fig. 12 reveal that scaling transformation stands out, achieving an unparalleled average tSuc of 58.0% at $s=0.8$. This performance is nearly 10% superior to that of the shear transformation (48.8%) and the perspective transformation (48.4%). Notably, the shear transformation shows substantial effectiveness at lower intensities; however, its efficiency markedly wanes at higher intensities. Furthermore, it is observed that employing color transformations in isolation offers marginal benefits for the attack. Yet, subsequent experimentation will demonstrate that when synergized as auxiliary augmentations to scaling, color transformations can significantly bolster targeted transferability.
Figure 12: Results by applying various basic transformations with incremental intensities. The scaling transformation exhibits a distinctive superior enhancement in targeted transferability.
4.2.3 Comparison among Scaling’s Variations
Figure 13: Our analysis uncovers significant untapped potential in existing scaling-related methods. We illustrate that prior achievements can be substantially enhanced by extending the scaling range and incorporating bidirectional scaling. Our tailored refinements to the S4ST-Base further amplify this effect.
We perform comparative evaluations of the (R)DI methods under both unidirectional (i.e., limited to either enlargement or reduction) and bidirectional conditions (i.e., integrating both enlargement and reduction), with the findings presented in Fig. 13. Specifically, under the hyperparameter configurations from prior TTA research, the DI and RDI methods attain an average tSuc of only 22.4% and 27.3%, respectively. Our research reveals that minor adjustments to their parameters can elevate these figures to 36.6% and 41.7%. By synergistically applying these methods in both the upscale and downscale directions, we obtain a significantly enhanced average tSuc of 56.8% at $r=2.4$, which, while slightly lower than the 57.2% achieved by the simple scaling transformation, is substantially surpassed by our S4ST-Base approach, which further boosts the outcome to 61.6%. This highlights the superiority of our method in both unlocking and augmenting the effectiveness of the basic scaling transformation. Fig. 14 delineates the time consumption of these methods, highlighting our approach’s computational efficiency. The incremental time investment required by S4ST-Base compared to the baseline method (at $r=1.0$) is minimal, and it markedly surpasses the competitors.
Figure 14: In addition to its effectiveness, our S4ST also exhibits superior efficiency. The additional time required by our S4ST is trivial, significantly outstripping the simple scaling and the integration of RDI [28] and DI [10] methods in terms of computational speed.
Figure 15: Further enhancements to S4ST-Base achieved by (a) S4ST-Block and (b) S4ST-Aug, both of which demonstrate significant improvements.
TABLE IV: Comparison between input transformations based on ResNet-50 as the surrogate model. strands for TTA-oriented methods. The best and the second best results are in bold and underlined, respectively.
Attack Venue tSuc Time MNv2 EN CNX INv3 INv4 IRv2 Xcep AvgC ViT Swin MaxViT Twins PiT TNT DeiT AvgT
TMI [8, 9] CVPR 2019 0.8 1.24 2.0 1.5 2.4 0.6 0.7 0.7 0.9 1.3 0.2 0.9 0.4 0.1 0.1 0.2 0.2 0.3
+Admix [45] ICCV 2021 2.1 6.27 5.9 4.7 6.3 0.9 1.6 2.4 2.5 3.5 0.3 2.0 1.0 0.5 1.1 0.6 0.2 0.8
+SI [44] ICLR 2020 4.4 6.11 8.5 9.6 8.4 5.7 5.3 6.3 7.0 7.3 1.1 2.4 1.5 1.2 2.5 1.3 0.6 1.5
+SSA [46] ECCV 2022 19.8 1.43 35.0 32.0 35.1 22.5 30.4 35.1 33.0 31.9 4.8 11.1 7.1 7.4 10.8 7.2 6.0 7.8
+DI [10] CVPR 2019 22.4 1.44 29.6 33.7 48.8 15.4 35.4 38.1 28.3 32.8 5.9 17.1 10.8 11.1 19.1 11.6 8.3 12.0
+RDI [28] ECCV 2020 27.3 1.24 37.1 42.9 47.4 25.8 39.8 45.6 36.9 39.4 11.0 21.7 13.2 13.6 22.3 15.0 10.2 15.3
+ODI [32] CVPR 2022 58.1 5.11 73.2 77.1 75.3 64.3 75.7 77.0 73.2 73.7 37.6 54.1 37.9 38.4 54.5 43.0 31.7 42.5
+SIA [11] ICCV 2023 58.3 1.55 82.5 81.3 84.3 51.1 74.2 76.2 72.0 74.5 30.8 58.5 42.6 39.3 53.2 40.7 29.8 42.1
+DeCoWA [13] AAAI 2024 58.4 2.78 74.9 78.0 77.4 56.7 74.4 77.7 73.7 73.3 40.6 58.4 36.1 36.6 55.1 44.1 33.3 43.5
+BSR [12] CVPR 2024 60.0 1.84 81.8 80.3 81.4 62.6 79.3 78.1 76.6 77.2 32.2 56.8 38.4 39.6 55.7 44.2 33.5 42.9
+T-Aug [27] ArXiv 2023 65.5 1.33 82.1 82.7 81.9 72.4 84.1 84.4 82.8 81.5 41.7 61.9 47.5 45.2 64.2 50.2 36.4 49.6
+H-Aug [27] ArXiv 2023 68.8 5.64 81.8 82.4 85.9 69.7 86.0 82.8 84.7 81.9 43.1 73.4 56.7 59.0 63.0 54.1 40.4 55.7
+S4ST Ours 83.0 1.45 93.5 93.9 93.9 85.9 93.0 92.9 93.1 92.3 63.0 86.3 74.8 73.8 80.1 75.1 63.2 73.8
[…]
Building upon the S4ST-Base framework, we delve into the benefits conferred by the proposed complementary augmentations (S4ST-Aug) and block-wise scaling (S4ST-Block). We maintain the probability of scaling transformations at $p_{\textit{r}}=0.9$ and progressively increase the number of image blocks $m$ and the probability of applying complementary augmentations $p_{\textit{Aug}}$. The outcomes are depicted in Fig. 15. For S4ST-Block, an escalation in the number of image blocks correlates with a marked enhancement in attack efficacy, attaining its zenith at $m=6\,(2\times 3)$. A further increase in blocks does not yield improved convergence within the configured iterations for the attack. In the case of S4ST-Aug, a higher frequency of applying complementary augmentations is linked with a steady amplification in performance. Additionally, the integration of five distinct types of complementary transformations, particularly when paired with scaling (for $r\geq 1.2$), results in substantially improved outcomes compared to their solitary application (at $r=1.0$). This emphasizes the soundness of our approach, which synergizes a primary transformation with complementary enhancements.
4.3 Comparison Against Existing Transformations
This subsection presents comparisons of the proposed S4ST with existing transformation methods across various settings.
4.3.1 Single Surrogate Attacks
Table IV reports the single surrogate model (ResNet-50) evaluation. Herein the transformed copies per iteration for SSA, SIA, BSR, and DeCoWa are limited to 1. We also set $m_{1}=1,m_{2}=5$ for Admix to ensure a fair comparison with SI. The results demonstrate that the proposed S4ST exhibits superior performance, not only achieving significantly higher average tSuc than existing transformation techniques, surpassing H-Aug by 14.2% and ODI by 24.9%, but also leading in terms of efficiency, with computational time only 25.7% of H-Aug and 28.4% of ODI. In fact, as can be seen in conjunction with Fig. 13, even the simple scaling transformation and our S4ST-Base are sufficient to achieve transferability on par with all methods except for T/H-Aug. And the further enhancements through block-wise scaling and complementary augmentations synergistically contribute to the final superior performance. This comparison is extended to alternative surrogate models (without further tuning the parameters), with results reported in Table V, and various attack iterations in Fig. 1. Our method consistently outperforms the others by clear margins.
TABLE V: Results by alternative surrogate models. strands for TTA-oriented methods. The best results are in bold.
Surrogate DenseNet-121 [69] VGG16-bn [70] ViT [60]
Attack tSuc Time tSuc Time tSuc Time
TMI [8, 9] 1.1 1.86 0.8 1.66 0.0 3.99
+ODI [32] 46.2 5.57 38.3 5.26 31.1 7.70
+SIA [11] 49.9 2.39 33.7 2.30 22.1 4.63
+DeCoWA [13] 43.6 3.89 46.5 3.66 14.6 8.30
+BSR [12] 47.9 2.03 32.1 1.95 27.6 4.27
+T-Aug [27] 49.3 2.01 48.1 1.75 27.2 4.06
+H-Aug [27] 59.3 4.51 48.3 7.00 - -
+S4ST (Ours) 74.6 2.15 65.6 1.83 64.8 4.05
4.3.2 Ensemble-based Attacks and Against Robust Victims
We further extend our comparison to scenarios involving the generation of AEs based on an ensemble of surrogate models and evaluations using robust victim models. For the selection of ensemble surrogate models, we distinguish between the commonly used self-ensemble in untargeted transfer attacks [46, 11, 13, 12], where the average gradient over multiple transformed copies is computed in a single iteration, and cross-model ensemble [8], which involves calculating the loss function using the average logits from different models to obtain gradients. In line with the evaluation protocols of previous research [15, 18], we incorporate robust victim models for evaluation, including Augmix [71], stylized training (SIN and SIN-IN) [72], adversarial training (AT) [2, 73], and Ensemble AT [3]. Table VI presents the results for tSuc and untargeted transfer success rate (uSuc) under the aforementioned settings.
TABLE VI: Results of attacking robust models (tSuc/uSuc) and ensemble-based attacks. stands for TTA-oriented methods. The best results are in bold. uSuc is the ratio (in %) of the false predictions caused by the targeted AE by the original correct predictions.
Attack Augmix [71] Stylized [72] Adversarial training (AT) [2] Ensemble AT [3] Avg. AvgC AvgT
[…]
Surrogate: ResNet-50
TMI [8, 9] 3.4/55.8 0.5/51.6 33.7/83.4 0.2/42.4 0.0/28.6 0.0/24.3 0.0/19.1 0.0/17.8 0.0/17.9 0.0/12.4 3.8/35.3 1.3 0.3
+ODI [32] 78.5/94.6 45.2/87.6 96.5/99.4 65.6/88.8 3.5/44.5 6.4/46.2 0.6/28.2 14.4/51.4 5.8/40.1 3.2/27.6 32.0/60.8 73.7 42.5
+SIA [11] 89.1/96.7 39.8/84.6 99.6/100.0 64.1/87.5 1.5/41.5 3.7/40.2 0.1/27.2 5.7/37.7 1.7/32.5 0.6/19.9 30.6/56.8 74.5 42.1
+DeCoWA [13] 82.8/95.7 45.2/87.9 97.7/99.8 72.2/91.7 4.0/48.5 8.2/49.0 0.7/30.2 16.8/58.2 8.8/49.0 4.0/34.1 34.0/64.4 73.3 43.5
+BSR [12] 86.5/96.1 48.0/87.0 98.9/99.9 69.6/89.9 2.8/44.9 5.5/46.9 0.3/25.7 11.0/48.7 3.8/38.5 2.1/27.6 32.9/60.5 77.2 42.9
+T-Aug [27] 87.4/97.1 44.2/88.6 97.6/99.8 71.8/90.0 3.7/46.8 7.0/47.2 0.3/28.8 19.5/57.1 8.0/46.1 4.6/33.4 34.4/63.5 81.5 49.6
+H-Aug [27] 80.7/94.5 34.8/81.0 93.4/97.8 62.0/86.7 2.8/43.9 5.6/44.4 0.3/27.9 13.0/49.6 5.4/39.0 2.5/26.6 30.1/59.1 81.9 55.7
+S4ST (Ours) 94.9/98.8 64.1/92.3 99.8/99.9 89.2/96.7 13.7/54.5 21.2/57.5 2.2/31.0 42.3/71.0 20.7/55.4 12.4/40.1 46.1/69.7 92.3 73.8
Surrogate: ResNet-50 $\times$ 10 (Self-Ensemble)
[…]
+ODI [32] 92.2/98.2 66.9/93.3 98.9/99.9 83.2/94.4 7.5/51.1 12.5/52.2 1.1/30.6 27.8/65.6 14.0/53.3 7.0/39.8 41.1/67.8 87.0 63.6
+SIA [11] 92.7/98.2 47.0/86.0 99.9/99.9 74.1/90.4 2.0/43.7 4.5/44.1 0.5/27.0 7.6/43.0 2.8/35.2 1.0/24.2 33.2/59.2 81.4 53.1
+DeCoWA [13] 88.1/96.9 56.4/93.4 97.8/99.8 78.0/92.8 4.7/47.8 8.5/50.2 0.8/31.0 18.7/59.9 9.3/49.6 4.4/36.2 36.7/65.8 78.4 52.8
+BSR [12] 91.3/97.8 54.4/88.7 99.5/99.9 74.0/91.2 2.8/44.6 6.3/47.9 0.5/27.1 11.6/52.0 5.6/42.6 2.6/31.3 34.9/62.3 82.6 51.5
+T-Aug [27] 92.6/98.3 56.2/91.1 98.6/99.8 81.1/93.5 5.7/51.1 10.8/51.0 0.5/29.5 24.6/62.4 11.5/50.4 6.5/37.0 38.8/66.4 87.3 62.6
+H-Aug [27] 92.8/98.2 58.4/90.5 98.1/99.8 81.4/94.7 8.0/49.4 13.5/51.7 0.6/30.0 29.8/65.0 14.2/51.7 6.9/37.6 40.4/66.9 90.4 73.3
+S4ST (Ours) 98.3/99.8 79.8/96.1 99.9/99.9 94.3/98.5 25.0/63.8 35.3/65.8 3.5/34.4 64.0/83.9 41.5/70.9 28.7/60.0 57.0/77.3 95.8 83.7
Surrogate: ResNet-50 + ResNet-152 + DenseNet-121 + VGG16-bn (Cross-Model Ensemble)
[…]
+ODI [32] 93.0/97.4 61.8/88.6 98.9/99.8 79.7/92.7 5.6/48.7 10.0/49.8 0.9/30.1 20.8/55.7 10.1/42.9 5.6/30.5 38.6/63.6 91.8 73.8
+SIA [11] 98.4/99.5 68.5/92.5 99.8/99.9 90.2/97.1 6.9/50.4 13.0/52.8 1.0/30.8 23.5/57.4 9.9/43.1 5.4/30.2 41.7/65.4 95.9 83.9
+DeCoWA [13] 97.7/99.0 77.0/96.2 99.4/99.9 93.4/98.0 13.4/57.7 21.7/59.8 2.4/35.4 51.9/79.7 32.5/65.6 17.7/49.0 50.7/74.0 95.2 82.2
+BSR [12] 98.1/99.6 73.0/94.2 99.8/99.9 89.8/96.8 7.3/52.0 15.4/55.9 1.0/31.3 35.0/67.1 14.0/51.3 8.4/37.1 44.2/68.5 95.4 78.7
+T-Aug [27] 97.6/99.4 67.7/93.9 99.7/99.9 90.3/97.4 10.5/52.5 16.5/57.4 1.4/33.3 48.8/75.4 24.3/56.9 14.7/44.8 47.2/71.1 96.3 83.1
+H-Aug [27] 97.1/99.1 62.9/91.2 99.4/99.9 88.1/96.0 8.6/52.2 14.8/54.6 1.1/31.5 39.4/69.0 17.1/52.6 9.1/39.6 43.8/68.6 96.5 85.3
+S4ST (Ours) 99.3/99.9 81.6/96.3 99.9/99.9 96.6/99.0 27.2/64.4 39.3/68.6 4.5/36.3 70.8/87.1 47.0/72.5 30.5/56.9 59.7/78.1 98.6 92.5
TABLE VII: Comparison with SOTA under the 10-Targets (all-source) setting. ##\## Data means the number of samples involved in method training. means the data is from the same dataset of victim models. means the victim models are involved for parameter tuning. The best results are in bold.
Attack Venue ##\## Data CNNs Vision transformers Robust mechanisms Avg.
[…]
DFA [14] CVPR 2020 50K‡◆ 25.3 1.6 15.0 16.3 13.8 0.7 3.1 1.6 1.7 3.9 19.9 7.8 56.4 0.3 0.3 11.2
TTP [15] ICCV 2021 50K‡◆ 54.7 38.4 56.6 57.9 60.6 30.0 35.6 24.4 28.6 11.4 68.7 27.8 84.1 2.9 5.4 39.1
M3D [18] CVPR 2023 50K‡◆ 84.0 65.9 83.5 85.9 84.4 60.8 67.7 62.0 62.2 41.0 88.0 62.9 94.6 5.2 9.2 63.8
C-GSP [37] ECCV 2022 1.2M‡◆ 57.2 20.5 47.1 48.5 46.7 22.6 26.5 19.7 22.1 9.8 50.7 17.4 72.1 1.4 1.8 30.9
IAA [19] ICLR 2022 5K‡◆ 63.2 30.7 55.2 58.7 54.6 15.9 33.2 24.6 27.5 17.0 81.6 42.7 98.7 1.4 2.4 40.5
DRA-DI-TMI [20] IEEE TIP 2022 1.2M‡◆ 58.3 69.2 58.9 67.9 61.7 69.1 55.9 63.1 55.6 46.6 86.5 90.2 94.3 73.4 71.8 68.2
LRS-DI-TMI [35] NeurIPS 2021 1.2M‡◆ 72.3 55.1 61.3 72.9 60.5 45.6 47.2 50.9 41.5 35.1 85.0 41.5 93.3 24.5 31.7 54.6
SASD-WS-DI-TMI [21] CVPR 2024 1.2M‡◆ 88.6 85.2 92.2 93.3 91.2 58.2 70.4 68.8 61.6 47.2 95.3 88.3 99.8 12.4 20.8 71.6
SU-DI-TMI [33] CVPR 2023 N/A 54.7 17.1 41.6 43.0 33.2 6.5 19.2 18.8 11.7 10.2 43.8 11.2 89.1 0.8 1.0 26.8
CFM-RDI-TMI [29] CVPR 2023 N/A 88.1 81.7 89.0 91.0 88.2 56.3 72.3 72.7 65.3 49.4 96.0 79.6 99.7 17.0 26.5 71.5
ODI-TMI [32] CVPR 2022 N/A 74.3 62.6 73.6 75.1 70.8 37.6 53.3 55.5 42.8 32.2 78.1 45.2 96.8 6.1 9.6 54.2
H-Aug-TMI [27] ArXiv 2023 N/A 86.3 69.0 84.2 82.8 82.5 42.9 72.3 65.2 57.2 40.9 80.8 36.4 92.7 5.5 8.7 60.5
S4ST-TMI Ours ${}^{\quad\!\!\!}$ 96.0 86.2 94.5 95.0 94.7 69.1 88.0 84.9 80.1 66.5 95.8 68.1 99.5 18.1 28.3 77.7
Firstly, the proposed S4ST transformation significantly outperforms existing methods across all scenarios. Moreover, upon closer examination, it is evident that S4ST benefits the most from increased computational effort and a larger number of surrogate models—whether it is averaging over multiple variants of a single surrogate model or computing gradients across models. Notably, it achieves performance close to that of the cross-model ensemble in the self-ensemble case, highlighting its potential for practical application. This is particularly relevant since obtaining multiple distinct surrogate models is generally challenging in the black-box scenario. Additionally, adversarially trained models with larger perturbation budgets exhibit significantly improved defensive capabilities, yet they remain susceptible to erroneous (untargeted) predictions induced by targeted AEs.
4.4 Comparison with SOTA
We further compare with SOTA approaches for both data-free and -reliant TTAs. Note that training universal perturbations [14] or single-target generators [15, 18] for all target labels specified in the ImageNet-Compatible dataset is computationally impractical. Therefore, we follow the 10-Targets (all-source) setting [15] that averages tSuc on 10 target labels—{Grey-Owl, Goose, Bulldog, Hippopotamus, Cannon, Fire-Truck, Model-T, Parachute, Snowmobile, Street sign} against the entire dataset. All methods are implemented targeting ResNet-50 as the surrogate model and the attack results on three types of models—CNNs, Vision Transformers, and robustly trained models—are reported in Table VII.
The proposed method achieves the best result with an average tSuc of 77.7%, outperforming the current best data-free method, CFM [29], and the best data-reliant method, SASD-WS [21], by 6.2% and 6.1%, respectively, even when they are already equipped with input transformation techniques. Moreover, this is achieved in a strict black-box manner. As summarized in the table, SASD-WS utilizes 1.2M samples of the ImageNet-1K training set for post-training of the surrogate model and CFM also requires the 1K evaluated data and access to black-box models for parameter tuning. Overall, our method offers the flexibility to handle any (sample, label) pair while achieving SOTA targeted transferability to date. It does not utilize any black-box models or their data during the method design and training process, further revealing the potential and risks of TTAs under strict black-box conditions.
TABLE VIII: Target labels and semantically similar alternatives of attacking real-world APIs.
APIs Street sign (T1) Goose (T2) Parachute (T3) Hippopotamus (T4) Cannon (T5)
Google Cloud Vision sign street name sign electronic signage canada goose ducks paragliding parachuting snout self-propelled artillery
Azure sign signage geese and swans canada goose paragliding parachuting hippo self-propelled artillery
Baidu {CJK}UTF8gbsn指路牌, {CJK}UTF8gbsn指示牌, {CJK}UTF8gbsn路标 {CJK}UTF8gbsn鹅, {CJK}UTF8gbsn加拿大鹅, {CJK}UTF8gbsn肉鹅, {CJK}UTF8gbsn野鹅, {CJK}UTF8gbsn天鹅 {CJK}UTF8gbsn降落伞, {CJK}UTF8gbsn动力伞, {CJK}UTF8gbsn滑翔伞 {CJK}UTF8gbsn河马 {CJK}UTF8gbsn火炮
TABLE IX: Results of attacking real-world APIs. The best results are in bold.
Attack ##\## Data Google Cloud Vision Azure Baidu Avg.
T1 T2 T3 T4 T5 Avg. T1 T2 T3 T4 T5 Avg. T1 T2 T3 T4 T5 Avg.
SASD-WS-DI-TMI [21] 1.2M 2 68 65 63 57 51.0 39 65 62 94 65 65.0 79 88 83 96 98 89.0 68.3
CFM-RDI-TMI [29] N/A 2 55 38 55 46 39.0 26 69 50 84 64 58.6 70 86 79 97 93 85.0 60.9
S4ST-TMI N/A 7 69 71 66 62 55.0 52 70 56 88 68 66.8 71 80 89 99 97 87.0 69.6
[…]
Figure 16: Real-world APIs-returned label list for targeted AEs generated by S4ST-TMI.
TABLE X: Results of attacking vision-language models. The best results are in bold.
Attack ##\## Data GLM-4V Flash [74] Gemini 1.5 Flash [75] Janus Pro 7B [76] Avg.
T1 T2 T3 T4 T5 Avg. T1 T2 T3 T4 T5 Avg. T1 T2 T3 T4 T5 Avg.
SASD-WS-DI-TMI [21] 1.2M 51 74 71 80 72 69.6 33 34 36 57 50 42.0 47 56 63 74 78 63.6 58.4
CFM-RDI-TMI [29] N/A 28 66 64 67 59 56.8 14 29 17 37 38 27.0 20 47 39 57 59 44.4 42.7
S4ST-TMI N/A 41 72 71 81 76 68.2 29 41 21 54 44 37.8 42 62 53 72 70 59.8 55.3
[…]
4.5 Transferring Against Real-World API
Evaluation against the commercially deployed models in the real world is performed to verify the practicality of our method. We follow the quantitative evaluation setting of [18] to perform transfer attacks on 5 labels over 100 images against 3 commercial APIs including the Google Cloud Vision (label detection) [^5] https://cloud.google.com/vision/docs/labels, Azure (image tagging) [^6] https://portal.vision.cognitive.azure.com/demo/generic-image-tagging, and Buidu (advanced image recognition) [^7] https://cloud.baidu.com/product/imagerecognition/general. tSuc here is measured if the target label appears in the predicted label list of the APIs. Additionally, to prevent the scenario where the label set of the remote model does not include the target label, we first test the clean samples of the target label and select the high-frequency predicted labels returned by the APIs as semantically similar labels, as clarified in Table VIII.
Table IX presents the quantitative results. Our method achieves superior performance over the SOTA data-reliant approach SASD-WS [21] by 1.3% (69.6% average tSuc) when using a single ResNet-50 surrogate model while employing ensemble surrogates further elevates the average tSuc to 91.7%. Fig. 16 provides several examples obtained from these APIs, demonstrating that targeted AEs can induce high-confidence erroneous predictions toward the target class. These results demonstrate that our S4ST-empowered attacks hold significant potential not only for evaluating model vulnerabilities under unknown training data distributions and data processing pipelines but also for safeguarding against malicious exploitation of AI models in applications such as CAPTCHA systems.
4.6 Evaluation on Vision-language Models
To further validate the scalability of our method to vision-language models, we maintain the evaluation setting of the previous subsection and design a prompt to assess attack effectiveness in this open-vocabulary scenario. Specifically, we instruct the models to describe the main object(s) in the targeted AEs and then select the best match from six shuffled options, which include the source label, target label, and four randomly sampled labels from the ImageNet-1K label set.
Three popular efficient vision-language models are tested including GLM-4V Flash [74], Gemini 1.5 Flash [75], and Janus Pro 7B [76], with results shown in Table X. In this scenario, our method achieves an average tSuc of 55.3% across the three models, which, although 3.1% lower than SASD-WS method, still demonstrates its competitiveness in enabling effective transfer attacks without relying on data or training. Furthermore, by leveraging multiple surrogate models, the performance can be significantly improved to 82.4%. Fig. 17 illustrates the responses of vision-language models to several targeted AEs, showing that our method effectively induces target label-related visual features and corresponding textual understanding.
Figure 17: Responses obtained from vision-language models for targeted AEs generated by S4ST-TMI. Source labels are in green and target labels are in red.
Figure 18: Targeted AEs, bolstered by S4ST, unveil the potential mechanism for enhancing targeted transferability through image transformations—by strengthening the manipulation of objects and textures within individual images to generate semantics that align with the intended target label.
4.7 Visual Inspections
Fig. 18 provides examples of targeted AEs. It demonstrates that our S4ST significantly enhances the infusion of the target label’s semantics into AEs, a process that is further augmented when attacking an ensemble of surrogate models. For instance, in the first set of examples, the bird crista is subtly altered to resemble the windows of a church loft; in the sixth set, the background among the flock of geese is imbued with the texture of a conch. These observations are consistent with previous findings on unbounded perturbations [24] and data-agnostic perturbations [15, 14], but they further underscore the manipulation of objects, textures, and backgrounds within individual images. More effective transformations can facilitate this process, revealing potential avenues for enhancing targeted transferability through transformations. However, this also reveals an intrinsic tension between stronger transferability and imperceptibility, which may provide a clue for distinguishing targeted AEs from untargeted ones and clean images, thereby aiding specializing defense strategies.
5 Conclusion
Two blind estimators, named self-alignment and self-transferability, have been proposed in this paper. They enabled analyses of the effectiveness and synergistic effect of various basic transformations in the strict black-box setting. The obtained findings have driven the design and tuning of a highly potent scaling-centered transformation, S4ST, which enjoys a significant increase in the transferability of data-free TTAs, all done without evaluation or training that involves any additional data and models. Compared with existing transformation techniques, our S4ST excels in efficiency, effectiveness, and adherence to black-box conditions. It also outperforms the SOTA solution of TTAs (no extra training versus retraining on 1.2M samples) in various settings including transferring against CNNs, vision transformers, robust models, commercial APIs, etc. We hope this will inspire the community to reconsider the data dependency of TTAs. Overall, this paper reveals the underdeveloped potential of data-free TTAs, and highlights the risks posed by black-box targeted attacks to real-world systems.
Aligning with advancements in adversarial machine learning, this research aims to deepen our understanding of adversarial vulnerabilities, thereby contributing to the development of trustworthy artificial intelligence systems [77]. The concept of targeted transferability is not only instrumental in enhancing privacy and intellectual property protection [78, 79] and enabling robust inference [2] but also critical for identifying potential risks in high-stakes applications such as autonomous driving [80, 81] and face recognition [82, 32]. While the potential for malicious exploitation of our findings cannot be entirely ruled out, we remain optimistic that this work will serve as a catalyst for stakeholders and the research community, inspiring the development of more robust defenses and ultimately driving positive societal impact.
Appendix A Details of Evaluated Basic Transformations
Our study investigates a range of basic transformations that are routinely utilized in training neural networks. To assess self-transferability and self-alignment, we strive to quantify the intensity of all transformations using a unified factor, denoted as $s\in[0,1]$. We establish the intensity factor set $\mathcal{S}$ as an arithmetic sequence comprising 100 equally spaced samples within the interval $[0,1]$. For transformations with multiple directions, we compile all variations at the same intensity to compute the average outcomes. For example, regarding rotation, we consider images rotated by an identical angle in both the clockwise and counterclockwise directions. Two special cases are translate and shear, for which we combine the variations in both horizontal and vertical directions and both the positive and negative angles. To encapsulate their multidirectional nature, we uniformly sample 100 grid points $(s_{1},s_{2})$ within the $[0,1]\times[0,1]$ space, yielding four variants at the same intensity level. We ensure a sufficiently broad range of variation for each transformation to secure more precise outcomes.
These transformations are executed using torchvision [^8] https://github.com/pytorch/vision/tree/main/torchvision/transforms, conforming to its parameter definitions. For an input $\bm{x}$ with dimensions $H\times W$, we succinctly describe the evaluation settings as follows.
  • Rotation returns two images respectively rotated by $\texttt{angle}=s\times 180^{\circ}$ (clockwise) and $\texttt{angle}=-s\times 180^{\circ}$ (counterclockwise) degrees.
  • Scaling returns two images respectively scaled to the size of $(1+1.5s)H\times(1+1.5s)W$ (enlargement) and $\frac{1}{(1+1.5s)}H\times\frac{1}{(1+1.5s)}W$ (reduction).
  • Shear returns four images respectively sheared by $(s_{1}\times 90^{\circ},s_{2}\times 90^{\circ})$, $(-s_{1}\times 90^{\circ},s_{2}\times 90^{\circ})$, $(s_{1}\times 90^{\circ},-s_{2}\times 90^{\circ})$, and $(-s_{1}\times 90^{\circ},-s_{2}\times 90^{\circ})$ degrees parallel to the x-axis and y-axis.
[…]
  • Flip returns two images respectively flipped vertically and horizontally.
  • Crop returns a square patch of $\bm{x}$ with an area of $(1-0.95s)\times H\times W$.
  • Translate returns four images respectively translated by $(s_{1}\times H,s_{2}\times W)$, $(-s_{1}\times H,s_{2}\times W)$, $(s_{1}\times H,-s_{2}\times W)$, and $(-s_{1}\times H,-s_{2}\times W)$ pixels along the x-axis and y-axis.
  • Solarize returns one image with pixels values $x_{i,j}\geq(1-s)$ been inverted to $1-x_{i,j}$.
  • Hue returns two images with the hue factor jittered to $-0.5s$ and $0.5s$, respectively.
  • Brightness/Contrast/Saturation returns two images with the brightness/contrast/saturation factor jittered to $(1+4s)$ and $\frac{1}{1+4s}$, respectively.